OpenStudy (anonymous):

A 1200 kg racecar is driven along a frictionless horizontal surface at a speed of 65 km/hr. If a horizontal cord brings the car to rest in a distance of 2.2 m, what is the elasticity constant (k) of this spring?

OpenStudy (shane_b):

Do you have any idea where to start?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

65km/hr=18.06m/s

OpenStudy (shane_b):

I'd start by converting the velocity to m/s. Then calculate the KE of the racecar using:\[KE=\frac{1}{2}mv^2\]

OpenStudy (shane_b):

Since the racecar comes to a rest you know that the change in KE will be the amount of work done on the car and: \[W=\frac{1}{2}kx^2\]

OpenStudy (anonymous):

1/2(1200kg)(18.06m/s)^2=195698.16J

OpenStudy (shane_b):

So:\[KE=\frac{1}{2}(1200kg)(18.06m/s)^2=195698J=Work\space done\]\[195698J=\frac{1}{2}k(2.2m^2)\]Now just solve for k.

OpenStudy (shane_b):

Small correction there...it should have been:\[195698J=\frac{1}{2}k(2.2m)^2\]

OpenStudy (anonymous):

Oh, because it's not the unit that's being squared.

OpenStudy (shane_b):

Yep

OpenStudy (shane_b):

Looking at that last equation, what do you get for k?

OpenStudy (anonymous):

k=80867.01

OpenStudy (anonymous):

What's that unit?

OpenStudy (shane_b):

It's a constant...no unit. You also have to consider significant figures in your calcs but that's how you solve it.

OpenStudy (anonymous):

Not overly concerned with significant figures in this case, but you make a good point. Thanks!

OpenStudy (shane_b):

no problem :)